OpenStudy (anonymous):

A ball is thrown vertically upward from the top of a 200 foot tower, with an initial velocity of 5 ft/sec. Its position function is s(t) = –16t2 + 5t + 200. What is its velocity in ft/sec when t = 3 seconds?

OpenStudy (anonymous):

Guys please Help me

OpenStudy (anonymous):

@radar could you please help me?

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

take the derivative and plug in 3 for t

OpenStudy (anonymous):

is that how you find the distance? @sourwing

OpenStudy (anonymous):

no. velocity

OpenStudy (anonymous):

WHT DO YOU MEAN BY DERIVATE?

OpenStudy (anonymous):

@sourwing ?

OpenStudy (anonymous):

Guys please help me

OpenStudy (anonymous):

@chrisdbest

OpenStudy (kash_thesmartguy):

Answer to a similar question - http://www.algebralab.org/Word/Word.aspx?file=Algebra_MaxMinProjectiles.xml

OpenStudy (anonymous):

they're similar but they talk about the maximums @Kash_TheSmartGuy

OpenStudy (kash_thesmartguy):

Oh ok.

OpenStudy (anonymous):

do you know someone that could help me? @Kash_TheSmartGuy

OpenStudy (kash_thesmartguy):

Probably @jim_thompson5910 or @campbell_st

OpenStudy (anonymous):

@jim_thompson5910 could you help me?

jimthompson5910 (jim_thompson5910):

this is for calculus class right?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

have you learned about derivatives? or differentiation?

OpenStudy (anonymous):

No I haven't, but my professor explained that it was related to the slope formula

jimthompson5910 (jim_thompson5910):

that's the average velocity, but they want the instantaneous velocity here

jimthompson5910 (jim_thompson5910):

it seems odd how he's asking about a topic you haven't learned about yet

OpenStudy (anonymous):

Wait I checked my notes isn't f(x) lim x-->0 f(a-h)-f(a) / h

jimthompson5910 (jim_thompson5910):

yes it is

jimthompson5910 (jim_thompson5910):

that's the limit definition of the derivative

OpenStudy (anonymous):

okay, but what could I first?

OpenStudy (anonymous):

Do I plug the 3 in the A right

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

what are f(3) and f(3+h) equal to

jimthompson5910 (jim_thompson5910):

btw that should be f(a+h) and not f(a-h) but I guess it doesn't matter if h goes to 0

OpenStudy (anonymous):

F(3) gives me 71 and F(a+h) gives me h^2+6h+9

jimthompson5910 (jim_thompson5910):

f(3) is correct but what you got for f(3+h) is not correct

OpenStudy (anonymous):

isn't -144-96h-16h^2

jimthompson5910 (jim_thompson5910):

f(t) = -16t^2+5t+200 f(3+h) = -16(3+h)^2+5(3+h)+200 f(3+h) = -16(9+6h + h^2)+5(3+h)+200 f(3+h) = -144-96h-16h^2+15+5h+200 f(3+h) = -16h^2-91h+71

jimthompson5910 (jim_thompson5910):

Use f(3) = 71 f(3+h) = -16h^2-91h+71 to find \[\Large \frac{f(3+h)-f(3)}{h}\]

OpenStudy (anonymous):

okay i see my mistake -91

jimthompson5910 (jim_thompson5910):

yep \[\Large \frac{f(3+h)-f(3)}{h}=\frac{-16h^2-91h+71-71}{h}\] \[\Large \frac{f(3+h)-f(3)}{h}=\frac{-16h^2-91h}{h}\] \[\Large \frac{f(3+h)-f(3)}{h}=\frac{h(-16h-91)}{h}\] \[\Large \frac{f(3+h)-f(3)}{h}=\frac{\cancel{h}(-16h-91)}{\cancel{h}}\] Then plug in h = 0 to get -16h-91 to turn into -91

jimthompson5910 (jim_thompson5910):

at that instant in time, the object is going -91 ft/sec (ie it's falling at 91 ft/sec)

OpenStudy (anonymous):

okay, OMG thank you very muchI will write this down about the instantaneous velocity formula. I really appreciate all your help, god will pay you in a way that you will feel blessed. thank you, I had to submit this today at 12 and I am ready, thank you so much

jimthompson5910 (jim_thompson5910):

you're welcome, I'm glad it helped out